OpenStudy (tvinsant):

PLEASE HELP ME. I WILL MEDAL AND FAN. 1. What is the equation, in standard form, of a parabola that contains the following points? Show your work. (–2, -20), (0, -4), (4, -20)

OpenStudy (anonymous):

do you know the standard formula of a parabola?

OpenStudy (tvinsant):

y=ax + bx + c

OpenStudy (tvinsant):

right?

OpenStudy (coconutjj):

do you know graph end behaviours

OpenStudy (anonymous):

close you forgot the square on the x in the first term.

OpenStudy (tvinsant):

no. i don't

OpenStudy (tvinsant):

oh yeah. i forgot. sorry.

OpenStudy (coconutjj):

how about quadratic formula

OpenStudy (anonymous):

so it should actually be \[ax^{2}+bx+c\]

OpenStudy (tvinsant):

I don't remember the quadratic formula.

OpenStudy (anonymous):

ok so it's saying that you have to have these three points so why not setup a system of equations. Do you know how to do so?

OpenStudy (tvinsant):

sorry no. could you walk me through it?

OpenStudy (anonymous):

sure so for each point take the equation and plug the point into the formula. So for the point (-2,-20) plug that into y= ax^2+bx+c. Show me how you do that.

OpenStudy (tvinsant):

so it would be 20=a(-2)^2+-2b+c

OpenStudy (anonymous):

minor mistake but it should be -20=a(-2)^2+-2b+c. ok so simplify that down.

OpenStudy (tvinsant):

oops. sorry again.

OpenStudy (anonymous):

no problem they are just careless mistakes. Just be careful during tests.

OpenStudy (gorv):

y=ax^2+bx+c

OpenStudy (tvinsant):

okay. -20=-4a+-2b+c ?

OpenStudy (gorv):

to find out a , b , c .....plug in the point in equation

OpenStudy (anonymous):

hmmm...-2^2 is what?

OpenStudy (gorv):

plug in each point and u will get 3 equation

OpenStudy (tvinsant):

omg. its 4….

OpenStudy (anonymous):

yup so your equation should be?

OpenStudy (tvinsant):

im so out of it….

OpenStudy (tvinsant):

-20=4a+-2b+c ?

OpenStudy (anonymous):

yes beautiful. Ok so now keep that aside and lets move onto the next point. (0,-4). What will the equation be, if we do the same thing?

OpenStudy (tvinsant):

-4=a(0)+b(0) +c

OpenStudy (anonymous):

simplify that down?

OpenStudy (tvinsant):

-4 = c ?

OpenStudy (anonymous):

Nice! Alright now for the last point of (4,-20)?

OpenStudy (tvinsant):

-20=16a+4b -4 -5 = 16a +4b

OpenStudy (anonymous):

so I agree with the first equation you posted:-20=16a+4b -4 but I don't know how you got to the next equation.

OpenStudy (tvinsant):

oh. was i supposed to simplify the first equation?

OpenStudy (anonymous):

yes! But you did so incorrectly lol!

OpenStudy (tvinsant):

lol. sorry!!! what should i have done?

OpenStudy (anonymous):

well tell me what you should do to simplify -20=16a+4b -4 down

OpenStudy (tvinsant):

divide -4 on both sides of the equation?

OpenStudy (anonymous):

mmm we could do that...but why can we first just add 4 to both sides?

OpenStudy (anonymous):

*can't

OpenStudy (tvinsant):

oh. to cancel it out?

OpenStudy (anonymous):

yes!

OpenStudy (tvinsant):

ok. so -16=16a+4b

OpenStudy (anonymous):

beautiful! Now we can divide 4 on each side like you wanted

OpenStudy (anonymous):

we could have done it at the beginning actually but this way takes less math.

OpenStudy (tvinsant):

lol okay. -4=16a + b

OpenStudy (anonymous):

well...16a/4 =?

OpenStudy (tvinsant):

4a

OpenStudy (anonymous):

so your equation should be?

OpenStudy (tvinsant):

-4=4a + b

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

so what are the three equations that we have so far?

OpenStudy (tvinsant):

-20=4a+-2b+c -4 = c -4=4a + b

OpenStudy (tvinsant):

oh wait...

OpenStudy (anonymous):

yes?

OpenStudy (tvinsant):

shouldn't i simplify the first one by plugging in the value of c?

OpenStudy (anonymous):

yes!

OpenStudy (tvinsant):

-16=4a-2b?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so now that we've used the second equation fully. What are the two equations that we have to work with?

OpenStudy (tvinsant):

-16=4a-2b -4=4a + b

OpenStudy (anonymous):

so you have an idea of how to simplify the two into 1 equation?

OpenStudy (tvinsant):

could you show me? i remember you have to somehow combine the two equations by canceling out one of the variables…i think..

OpenStudy (anonymous):

haha that didn't come out at all...but sure, why not cancel out the equations but substracting one from another like this:  do you know how to do so?

OpenStudy (tvinsant):

like multiply the bottom equation by -1...

OpenStudy (anonymous):

or we could do it that way of isolating one variable and substituting, it's up to you,

OpenStudy (tvinsant):

which one is easier?

OpenStudy (tvinsant):

how would you do that>

OpenStudy (anonymous):

the first way. as for how you do it. Just simply subtract like you would normally do it so what's -16 - -4? Do you see how that works?

OpenStudy (tvinsant):

okay. so then it would be: -12=-3b?

OpenStudy (anonymous):

beautiful!

OpenStudy (anonymous):

simplify?

OpenStudy (tvinsant):

-4 = b?

OpenStudy (anonymous):

well...-12/-3 is?

OpenStudy (tvinsant):

4

OpenStudy (anonymous):

so?

OpenStudy (tvinsant):

4 = b

OpenStudy (anonymous):

yes now we can solve for a can't we?

OpenStudy (tvinsant):

yeah. so -2=a

OpenStudy (anonymous):

yes! so what's our entire equation?

OpenStudy (tvinsant):

y = -2x^2 + 4 + -4

OpenStudy (anonymous):

where's the x for the second term?

OpenStudy (tvinsant):

oops…. y = -2x^2 + 4x + -4

OpenStudy (anonymous):

nice! and there you go! Your answer!

OpenStudy (tvinsant):

Thank you so much! :))))

OpenStudy (anonymous):

no problem! You did the work!

OpenStudy (tvinsant):

lol. only cos' i was guided by such a smart teacher.

OpenStudy (anonymous):

aww thanks...but not really lol.